OpenStudy (anonymous):

Estimate by first rounding each number to the nearest integer. -5 1/8 + 2 5/6. A. –3 B. –2 C. 2 D. 7 A. –3 B. –2 C. 2 D. 7

OpenStudy (anonymous):

@iGreen @e.mccormick @Abhisar

OpenStudy (anonymous):

@shinalcantara

OpenStudy (shinalcantara):

-5 1/8 + 2 5/6 = -41/8 + 17/6 getting the LCD you'll have -41/8 + 17/6 = -123/24 + 68/24 = -55/24 = -2 7/24